A man jumps horizontally from the third floor of a burning building to a rescue trampoline below. His initial speed is 3.00 m/s. If the third floor window is 11.9 m above the trampoline, how much time (in seconds) will elapse between when he jumps and when he lands on the trampoline?
Answer #1

Given,

The initial speed, u = 3 m/s

The height of window, h = 11.9m

By using the equation of motion,

v2 - u2 = 2 g h

v2 - (3)2 = 2 * 9.8 * 11.9

v2 = 242.24

v = 15.56 m/s

We know,

v = u + g t

t = (v - u) / g

t = (15.56 - 3) / 9.8

t = 1.282 sec